On September 1, Shawn Dahl established Whitewater Rentals, a canoe and kayak rental business. The following transactions occurred in the month of September and affected the following accounts: Cash Accounts Payable Accounts Receivable Shawn Dahl, Capital Office Equipment Revenue Canoe and Kayak Equipment Expenses Transactions Shawn Dahl invested $45,100 in cash to open the business Paid $12,800 in cash for the purchase of kayak and canoe equipment Paid $1,100 in cash for rent expense Purchased additional kayak and canoe equipment for $3,850 on credit Received $3,950 in cash for kayak rentals Rented canoes and kayaks for $1,250 on account Purchased office equipment for $130 in cash Received $825 in cash from credit clients Shawn Dahl withdrew $1,550 in cash for personal expenses Based on the information above, complete the following accounting equation.

Assets _=Liabilities_ + Owner's Equity _
Answer #1
Assets = Liabilities + Owner's Equity
Cash + Accounts Receivable + Office Equipment kayak and canoe equipment = Accounts Payable + Shawn Dahl, Capital + Revenues - Expenses - Shawn Dahl, Drawings
1 45,100 + + = + 45,100 + - -
2 -12,800 + + 12,800 = + + - -
3 -1,100 + + = + + - 1,100 -
4 + + 3,850 = 3,850 + + - -
5 3,950 + + = + + 3,950 - -
6 + 1,250 + = + + 1,250 - -
7 -130 + + 130 = + + - -
8 825 + -825 + = + + - -
9 -1,550 + + = + + - - 1,550

Kindly comment if you need further assistance. Thanks‼!